1. Assume the following balance sheet for a commercial bank: Assets Liabilities Reserves 100 Demand Deposits 1000 Government Bonds 400 Time Deposits 500 Mortgages 1000 Commercial Paper 400 Loans 500 Capital 100 Remember, for the balance sheet to balance, assets=liabilities + capital (or shareholder equity) The reserve requirement is 10% of demand deposits.

a. Suppose the bank is required to keep 15% of its risk-weighted assets in the form of capital. The risk weights are 0 for reserves and government securities, 20% for mortgages and 100% for loans. Does the bank have enough capital?

b. Suppose the bank earns a 2% return on its holdings of government securities, a 4% return on its mortgages and a 9% return on its loans. What is the bank’s return on equity?

c. Suppose the bank wants to increase the amount of its capital. Explain one way that it could do this.

d. Suppose the value of the mortgages falls to 900 because some of the borrower’s default on their payments. Discuss the effect this will have on the balance sheet and solvency of the bank.

e. Discuss options the bank can pursue if they run into liquidity (but not solvency) issues. Which options would the bank prefer and why?
Answer #1
Q a)
Particulars Amount
Liabilities
Capital (Balancing Fig.) 700
Reserves 100
Demand Deposits 1000
Time Deposits 500
Total 2300
Assets
Mortgages 1000
Government Bonds 400
Commercial Papers 400
Loans 500
Total 2300
15% of risk weighted assets:
Particulars Amount Risk Risk Weighted Asset
(a) (b) (c ) (d = b*c)
Government Bonds 400 0% 0
Commercial Papers 400 0% 0
Mortgages 1000 20% 200
Loans 500 100% 500
Total 2300 700
Therefore, risk weighted assets is equal to 700 and the bank has kept enough balance in its capital account.
Q b)
Particulars Amount Return (%) Return
(a) (b) (c ) (d = b*c)
Government Bonds 400 2% 8
Commercial Papers 400 2% 8
Mortgages 1000 4% 40
Loans 500 9% 45
Total 101
Total Equity of the bank = capital + reserves
= 700 + 100
= 800
Return on equity
= 101/800
12.63%
Q c)
If the bank wants to increase its capital, one way to increase the capital is issuance of shares. If laws governing banking system in the country allow the bank to issue shares, then, the bank shall issue shares to increase its capital
Q d)
If the value of mortgage decreases to 900, it will result in overall reduction in capital and also reduction of Asset balance.
Revised balance sheet will be displayed as follows:
Particulars Amount
Liabilities
Capital (Balancing Fig.) 600
Reserves 100
Demand Deposits 1000
Time Deposits 500
Total 2200
Assets
Mortgages 900
Government Bonds 400
Commercial Papers 400
Loans 500
Total 2200
Q e)
If the bank runs into liquidity issues and not solvency issues, it shall issue short term securities with maturities of 14 days, this will lead to short term cash inflow and further, it will not require the bank to pay much interest.